OpenStudy (anonymous):

The formula for the perimeter of a rectangle is P = 2l + 2w. Find l, if P = 60 and w = 14. Answers l = 16 l = 32 l = 28 l = 30 Can Anyone Help Me Out !

OpenStudy (anonymous):

\[60=2*14+2*w\]

OpenStudy (anonymous):

60 = 2l+ 2(14) 60 = 2l + 28 -28 = -28 32 = 2l ------ 2 L= 16

OpenStudy (anonymous):

\[60=2*l+2*14\] \[60=2*l+28\] \[60-28=2*l\] \[32=2*l\] \[l=16\]